what is the point of intersection for each pair of straight line by solving simultaneous equation
y= 2/3x +3
y= 4/3x +1
can anyone help me?

Answer by dfrazzetto(283)

y= 2/3x +3
y= 4/3x +1
y - 2/3x = 3
y - 4/3x = 1
Solved by pluggable solver: Linear System solver (using determinant)
Solve:


Any system of equations:


has solution

or



(y=5.00000000000001, x=3.00000000000002}


x=3, y=5